# Astrophysical Consequences of Black Holes in Asymptotically Safe Quantum Gravity: Shadows and Strong Gravitational Lensing

### v2

@@ II Black hole solutions in asymptotically safe quantum gravity
Black holes in asymptotically safe quantum gravity are described by solutions derived from the quantum effective action and the associated quantum equations of motion. These solutions are formulated using multi-graviton correlation functions within the asymptotically safe quantum gravity framework and have been recently explored numerically by Pawlowski and Tränkle [101]. The effective action is given by
@@
Here, the functions $f_{R^{2}}(\Delta)$ and $f_{R_{\mu\nu}^{2}}(\Delta)$ correspond to two-point correlation functions of the curvature scalar and Ricci tensor, respectively. These functions, often referred to as form factors, capture generalized dispersion effects. Studies of such form factors have been presented in works such as [102] and [103, 104, 105, 106]. For a comprehensive review of recent developments, see [107]. To derive black hole solutions within the framework of modified asymptotically safe gravity in quantum theory, one must solve the field equations obtained from the effective action $\Gamma[\bar{g}]$ given by Eq. (1) [101].
The static, spherically symmetric metric in asymptotically safe quantum gravity can be expressed as [101]:
@@
The two free parameters, $S_{0}$ and $S_{2}$, govern the strength of the exponentially decaying Yukawa corrections. The masses $m_{0}$ and $m_{2}$ correspond to the spin-0 and spin-2 modes, respectively, and are linked to the couplings of various curvature terms. For the initial conditions $(S_{0},S_{2})=(0,0)$, the asymptotically safe quantum gravity black hole metric (2) reduces to the classical Schwarzschild metric. The event horizon, located at $r_{h}$, is determined by the solution to the equation $g(r_{h})=0$.
III Shadow of a black hole in asymptotically safe quantum gravity
In this section, we initiate our investigation of the black hole shadow within the asymptotically safe quantum gravity framework, focusing on the motion of photons and the resulting shadow. Our analysis begins by examining the trajectories of light in the spacetime of an asymptotically safe quantum gravity black hole. Furthermore, we aim to constrain the black hole parameters associated with this asymptotically safe quantum gravity quantum gravity model. Now using Eq. (3) we can introduce the function $\gamma(r)$ (it was recently introduced in Refs. [108, 109, 110]):
@@ III Shadow of a black hole in asymptotically safe quantum gravity
The photon orbit radii $r_{ph}$ in asymptotically safe quantum gravity can be determined numerically using this equation.
When the parameters $S_{0}=S_{2}=0$, the solution yields $r_{ph}=3M$ [108, 110], corresponding to the photon orbit radius in the Schwarzschild spacetime. We now analyze the shadow of the black hole within the asymptotically safe quantum gravity framework.
@@
where $\alpha_{sh}$ denotes the angular radius of the black hole shadow, and $r_{obs}$ is the observer’s distance, assumed to be very large but finite. For instance, the observer distance for Sgr A* is $r_{obs}=D\simeq 8.3\,\text{kpc}$ [28], while for M87*, it is $r_{obs}=D\simeq 16.8\,\text{Mpc}$ [5].
@@
This equation allows us to examine how the shadow radius of the black hole depends on the parameters of asymptotically safe quantum gravity.
The dependence of the black hole shadow on the parameters $S_{0},~{}S_{2},~{}m_{0}$, and $m_{2}$ is illustrated in Fig. 1. It is evident that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the size of the black hole shadow decreases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, increases the size of the black hole shadow.
Figure 1: The shadow silhouette of the black hole in asymptotically safe quantum gravity is presented in the X-Y plane ($R_{sh}=\sqrt{X^{2}+Y^{2}}$ ), with black hole parameters $S_{0}$,$S_{2}$,$m_{0}$ and $m_{2}$ varied independently, while the other parameters are kept fixed.
III.1 Constraints the black hole parameters with the observational data from M87* and Sgr A*
Now, we aim to determine the possible ranges of the free parameters of the black hole in asymptotically safe quantum gravity by utilizing observational data from the supermassive black holes M87* and Sgr A*. Specifically, we will explore the theoretical constraints on the black hole parameters $S_{0}$, $S_{2}$, $m_{0}$, and $m_{2}$ pairwise while keeping the remaining two parameters fixed.
To achieve this, we use the shadow data provided by the Event Horizon Telescope (EHT) project. The objective is to constrain the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ by fitting them against the observed shadow sizes of M87* and Sgr A* within the framework of asymptotically safe quantum gravity. This approach will allow us to impose meaningful constraints on these parameters using real observational data.
The angular diameter of the shadow, the distance from the Sun to the black hole, and the mass of the black hole at the center of the galaxy M87 are given as $\Omega_{\text{M87*}}=42\pm 3\>\mu\text{as}$, $D=16.8\pm 0.8\>\text{Mpc}$, and $M_{\text{M87*}}=(6.5\pm 0.7)\times 10^{9}\>M_{\odot}$, respectively [112].
For Sgr A*, the most recent data obtained by the EHT project are $\Omega_{\text{Sgr A*}}=51.8\pm 2.3\>\mu\text{as}$, $D=8277\pm 9\pm 33\>\text{pc}$, and $M_{\text{Sgr A*}}=4.297\pm 0.013\times 10^{6}\>M_{\odot}$ (from VLTI observations) [28].
Using these observational data, we can estimate the diameter of the shadow cast by the black hole per unit mass, using the following expression [111, 113, 114]:
@@ III.1 Constraints the black hole parameters with the observational data from M87* and Sgr A*
Now we can obtain the diameter of the shadow from the expression $d_{\text{sh}}^{\text{theo}}=2R_{\text{sh}}$. Thus, the diameter of the BH shadow image is $d^{\text{M87*}}_{\text{sh}}=(11\pm 1.5)M$ for M87* and $d^{\text{Sgr A*}}_{\text{sh}}=(9.5\pm 1.4)M$ for Sgr A*. The angular diameters of the shadows of the black hole in asymptotically safe quantum gravity, obtained using the M87* and Sgr A* data, are depicted in Fig. 2, respectively.
@@
Table 1: Estimated ranges of the asymptotically safe quantum gravity black hole parameters $S_{0},~{}S_{2}$ and $m_{0},~{}m_{2}$ from the known shadow observables $\Omega(\mu as)$ for $M87^{*}$ and $Sgr~{}A^{*}$.
@@
Figure 2: The angular diameter $\Omega(\mu as)$ of the black hole shadow is presented as a function of the parameters $S_{0},~{}S_{2}$ and $m_{0},~{}m_{2}$ for the M87* (panels (a) and (c)) and Sgr A* (panels (b) and (d)) supermassive black holes. The enclosed regions satisfy the 1-$\sigma$ shadow bounds for M87* and Sgr A*, respectively. In panel (a), the green and blue lines correspond to $\Omega=39.1\,\mu\text{as}$ and $\Omega=39.5\,\mu\text{as}$, respectively; in panel (b), they correspond to $\Omega=51.4\,\mu\text{as}$ and $\Omega=51.8\,\mu\text{as}$, respectively; in panel (c), to $\Omega=39.4\,\mu\text{as}$ and $\Omega=39.3\,\mu\text{as}$, respectively; and in panel (d), to $\Omega=52.2\,\mu\text{as}$ and $\Omega=52\,\mu\text{as}$, respectively.
Constraints from M87*
@@
The shadow’s angular diameter ($\Omega$) for the M87* black hole, averaged within the range of $32$ to $45\>\mu\text{as}$ with a $1\sigma$ region from EHT observations, imposes constraints on the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ of the black hole in asymptotically safe quantum gravity, as shown in Figs. 2(a) and 2(c).
@@
Here, the shadow’s angular diameter ($\Omega$) is expressed as a function of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$. In Fig. 2(a), the green and blue lines correspond to $\Omega=39.1\>\mu\text{as}$ and $\Omega=39.5\>\mu\text{as}$, respectively. Similarly, in Fig. 2(c), the green and blue lines correspond to $\Omega=39.4\>\mu\text{as}$ and $\Omega=39.3\>\mu\text{as}$, respectively. The blue and green curves in Figs. 2(a) and 2(c) align with the shadow size of M87* observed by the EHT.
@@
The constraints on the parameter pair $(S_{0},S_{2})$ are given by (see Table 1 ): $0\leq S_{2}/2M\leq 0.09,$ and $0\leq S_{2}/2M\leq 0.09,$ corresponding to $\Omega=39.1\>\mu\text{as}$. Similarly, the constraints on the parameter pair $(m_{0},m_{2})$ are: $0\leq 2Mm_{0}\leq 0.78,$ and $0\leq 2Mm_{0}\leq 0.78$,corresponding to $\Omega=39.4\>\mu\text{as}$.
@@
Thus, M87* can be modeled within this constrained parameter space as a black hole in asymptotically safe quantum gravity. The numerous possible parameter points within the parameter spaces $(S_{0},S_{2})$ and $(m_{0},m_{2})$ indicate that the compatibility of the black hole model in asymptotically safe quantum gravity with M87* observations makes it a strong candidate for describing astrophysical black holes.
@@
The shadow’s angular diameter ($\Omega$) for the Sgr A* black hole, averaged within the range of $49.5$ to $54.1\>\mu\text{as}$ with a $1\sigma$ region from EHT observations, imposes constraints on the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ of the black hole in asymptotically safe quantum gravity, as shown in Figs. 2(b) and 2(d).
Here, the shadow’s angular diameter ($\Omega$) is expressed as a function of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see in Fig. 2). In Fig. 2(b), the green and blue lines correspond to $\Omega=51.4\>\mu\text{as}$ and $\Omega=51.8\>\mu\text{as}$, respectively. Similarly, in Fig. 2(d), the green and blue lines correspond to $\Omega=52.2\>\mu\text{as}$ and $\Omega=52\>\mu\text{as}$, respectively. The blue and green curves in Figs. 2(b) and 2(d) align with the shadow size of Sgr A* observed by the EHT.
The constraints on the parameter pair $(S_{0},S_{2})$ are given by (See Table 1 ): $0\leq S_{0}/2M\leq 0.11,$ and $0\leq S_{2}/2M\leq 0.11,$ corresponding to $\Omega=51.4\>\mu\text{as}$. Similarly, the constraints on the parameter pair $(m_{0},m_{2})$ are: $0\leq 2Mm_{0}\leq 0.78$ and $0\leq 2Mm_{2}\leq 0.78$, corresponding to $\Omega=52.2\>\mu\text{as}$.
Thus, Sgr A* can be modeled as a black hole in asymptotically safe quantum gravity within this constrained parameter space. The numerous possible parameter points within the parameter spaces $(S_{0},S_{2})$ and $(m_{0},m_{2})$ indicate that the compatibility of the black hole model in asymptotically safe quantum gravity with Sgr A* observations makes it a strong candidate for describing astrophysical black holes.

### v3

@@ II Black hole-like solutions in quadratic gravity
To explore how black holes are modified within the framework of quantum theory, one can investigate them in the context of quadratic gravity. Black holes in quadratic gravity are described by solutions derived from the quantum effective action and the associated quantum equations of motion. These solutions are formulated using multi-graviton correlation functions within quadratic gravity framework and have been recently explored numerically by Pawlowski and Tränkle [102]. The effective action is given by
@@
Here, the functions $f_{R^{2}}(\Delta)$ and $f_{R_{\mu\nu}^{2}}(\Delta)$ correspond to two-point correlation functions of the curvature scalar and Ricci tensor, respectively. These functions, often referred to as form factors, capture generalized dispersion effects. Studies of such form factors have been presented in works such as [103] and [104, 105, 106, 107]. For a comprehensive review of recent developments, see [108]. To derive black hole-like solutions within the framework in quadratic gravity, one must solve the field equations obtained from the effective action $\Gamma[\bar{g}]$ given by Eq. (1) [102, 21].
The static, spherically symmetric metric in quadratic gravity can be expressed as [21, 102]:
@@
The two free parameters $S_{0}$ and $S_{2}$ govern the strength of the exponentially decaying Yukawa corrections. The masses $m_{0}$ and $m_{2}$ correspond to the spin-0 and spin-2 modes, respectively and are linked to the couplings of various curvature terms. For the initial conditions $(S_{0},S_{2})=(0,0)$, the asymptotically safe quantum gravity black hole metric Eq.(2) reduces to the classical Schwarzschild metric. The event horizon, located at $r_{h}$, is determined by the solution to the equation $g(r_{h})=0$.
III Shadow of a black hole in quadratic gravity
In this section, we begin our investigation of the black hole shadow within the quadratic gravity framework, focusing on the motion of photons and the resulting shadow. Our analysis begins by examining the trajectories of light in the space-time of a quadratic gravity black hole. Furthermore, we aim to constrain the black hole parameters associated with this quadratic gravity model. Now using Eq. (3) one can consider the function $\gamma(r)$ [109, 110, 111]:
@@ III Shadow of a black hole in quadratic gravity
The photon orbit radii $r_{ph}$ in quadratic gravity can be determined numerically using this equation.
When the parameters are $S_{0}=S_{2}=0$, the solution yields $r_{ph}=3M$ [109, 111], corresponding to the radius of the photon orbit in the Schwarzschild spacetime.
Evaluating Eq. (6) for the linearized solution, as outlined in the literature [24], the photon sphere radius $r^{L}_{ph}$ in the linear approximation for the metric (2) is given by:
$r^{L}_{ph}=3M+\frac{3}{2}S_{2}(m_{2}M-1)e^{-3m_{2}M}+\frac{3}{2}S_{0}(m_{0}M-1)e^{-3m_{0}M}.$ (7)
Figure 1: Behaviour of event horizon radius $r_{h}/2M$, photon sphere radius for standard Schwarzschild black hole $r_{sh}/2M=1.5$, standard photon sphere radius $r_{ph}/2M$ and new photon sphere radius $r^{L}_{ph}/2M$ with the parameters $S_{0}/2M$ (left panel) and $2Mm_{0}$ (right panel), keeping the other parameters fixed.
$S_{0}/2M$ $2Mm_{0}$ $S_{2}/2M$ $2Mm_{2}$ $r_{h}/2M$ $r_{ph}/2M$ $r^{L}_{ph}/2M$ $R_{sh}/2M$ $R^{L}_{sh}/2M$
 0     0 0       1 1.5 1.5 2.598 2.598
$0.3$ 1.042 1.289 1.406 2.287 2.31
$0.1$ $0.6$ 1.047 1.32 1.426 2.349 2.367
$0.9$ 1.052 1.343 1.436 2.39 2.404
1.1 1.056 1.356 1.44 2.41 2.422
$0.3$ 0.995 1.164 1.366 2.09 2.154
$0.5$ $0.2$ $0.6$ 1.003 1.223 1.404 2.214 2.265
$0.9$ 1.014 1.272 1.425 2.301 2.338
1.1 1.022 1.299 1.434 2.343 2.372
$0.3$ 0.947 1.031 1.325 1.877 2.006
$0.3$ $0.6$ 0.959 1.116 1.383 2.062 2.166
$0.9$ 0.974 1.192 1.415 2.201 2.273
1.1 985 1.234 1.427 2.268 2.323
$0.1$
$0.3$ 1.025 1.329 1.443 2.357 2.378
$0.1$ $0.6$ 1.029 1.36 1.462 2.419 2.436
$0.9$ 1.034 1384 1.473 2.461 2.473
1.1 1.038 1.397 1.476 2.48 2.49
$0.3$ 0.978 1.203 1.402 2.162 2.222
$1$ $0.2$ $0.6$ 0.986 1.264 1.441 2.288 2.335
$0.9$ 0.996 1.314 1.462 2.375 2.408
1.1 1.004 1.341 1.47 2.417 2.442
$0.3$ 0.931 1.069 1.361 1.951 2.074
$0.3$ $0.6$ 0.942 1.158 1.419 2.14 2.237
$0.9$ 0.957 1.126 1.451 2.279 2.344
1.1 0.968 1.278 1.464 2.346 2.395
Table 1: Estimation of the event horizon radius $r_{h}/2M$, standard photon sphere radius $r_{ph}/2M$, new photon sphere radius $r^{L}_{ph}/2M$, black hole shadow radius $R_{sh}/2M$ corresponding to $r_{ph}/2M$ and black hole shadow radius $R^{L}_{sh}/2M$ corresponding to $r^{L}_{ph}/2M$ for different values of the parameters $S_{0}=0,0.1$, $S_{2}=0,0.1,0.2,0.3$, $m_{0}=0.5,1$ and $m_{2}=0.3,0.6,0.9,1.1$ respectively.
From the perspective of quadratic gravity, the geometry described by Eqs. (2)-(4) represents the gravitational field in the linear approximation. However, this framework breaks down in the strong gravity regime, where the event horizon becomes unstable due to the inclusion of exponentially small Yukawa corrections. Additionally, non-zero values of $S_{0}$ and $S_{2}$ can give rise to wormholes and naked singularities, leading to significant deviations from standard black hole geometry. These deviations necessitate careful consideration when applying the standard formulas for shadow imaging, as new photon paths may emerge that are otherwise blocked by the event horizon in conventional scenarios. To address these complexities, we have calculated the event horizon radius $r_{h}$, the standard photon sphere radius $r_{\text{ph}}$ and the new photon sphere radius $r_{\text{ph}}^{L}$, obtained through a linearized solution approximation. Our analysis shows that $r_{h}<r_{\text{ph}}<r_{\text{ph}}^{L}$ (see Table 1) based on our specific choice of parameter set, indicating that the photon sphere and its modifications, such as the new photon sphere obtained through the linearized solution approximation, play a crucial role in determining the shadow size. From Fig. 1, it is observed that for an arbitrarily chosen parameter set, the relation $r_{\text{ph}}<r_{\text{ph}}^{L}$ is maintained. However, the standard photon sphere radius can sometimes be smaller than the event horizon radius depending on the parameter choice, implying that the standard photon sphere may be blocked by the event horizon in certain cases. Otherwise, when both the standard photon sphere and the new photon sphere are present, this dual-photon-sphere configuration significantly influences the formation of the black hole shadow.
We now analyze the shadow of the black hole within the quadratic gravity framework.
@@
where $\alpha_{sh}$ denotes the angular radius of the black hole shadow and $r_{obs}$ is the observer’s distance, assumed to be very large but finite. For instance, the observer distance for $SgrA^{*}$ is $r_{obs}=D\simeq 8.3\,\text{kpc}$ [31], while for $M87^{*}$, it is $r_{obs}=D\simeq 16.8\,\text{Mpc}$ [5].
@@
This equation allows us to examine how the shadow radius of the black hole depends on the parameters of quadratic gravity.
The dependence of the black hole shadow on the parameters $S_{0},~{}S_{2},~{}m_{0}$ and $m_{2}$ is illustrated in Fig. 2. It is evident that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while the other parameters remain fixed, the size of the black hole shadow decreases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, increases the size of the black hole shadow. We find that the shadow size calculated using the standard photon sphere radius $r_{\text{ph}}$ is smaller than that obtained using the new photon sphere radius $r_{\text{ph}}^{L}$. These results underscore the importance of incorporating the effects of the new photon sphere radius when investigating shadow imaging in quadratic gravity, particularly in the strong gravity regime.
Figure 2: The shadow silhouette of the black hole in quadratic gravity is presented in the X-Y plane ($R_{sh}=\sqrt{X^{2}+Y^{2}}$ ), with black hole parameters $S_{0}$,$S_{2}$,$m_{0}$ and $m_{2}$ varied independently, while the other parameters are kept fixed. The dotted circular ring correspondence to shadow radius $R_{sh}/2M$ corresponding to $r_{ph}/2M$ and the solid circular ring correspondence to shadow radius $R^{L}_{sh}/2M$ corresponding to new photon sphere radius $r^{L}_{ph}/2M$ $r^{L}_{ph}/2M$.
III.1 Constraints the black hole parameters with the EHT observational data from $M87^{*}$ and $SgrA^{*}$
Here, we aim to constrain the possible values of the free parameters of black holes in quadratic gravity by analyzing observational data from supermassive black holes $M87^{*}$ and $SgrA^{*}$. Specifically, we will explore the theoretical constraints on the black hole parameters $S_{0}$, $S_{2}$, $m_{0}$ and $m_{2}$ pairwise while keeping the remaining two parameters fixed.
To achieve this, we use the shadow data provided by the Event Horizon Telescope (EHT) project. The objective is to constrain the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ by fitting them against the observed shadow sizes of $M87^{*}$ and $SgrA^{*}$ within the framework of quadratic gravity. This approach will allow us to impose meaningful constraints on these parameters using real observational data.
The angular diameter of the shadow, the distance from the Sun to the black hole, and the mass of the black hole at the center of the galaxy M87 are measured to be $\Omega_{\text{M87*}}=42\pm 3\,\mu as$, $D=16.8\pm 0.8\,\text{Mpc}$, and $M_{\text{M87*}}=(6.5\pm 0.7)\times 10^{9}\,M_{\odot}$, respectively [113].
For $SgrA^{*}$, the most recent data obtained by the EHT project are $\Omega_{\text{Sgr A*}}=51.8\pm 2.3\>\mu as$, $D=8277\pm 9\pm 33\>\text{pc}$ and $M_{\text{Sgr A*}}=4.297\pm 0.013\times 10^{6}\>M_{\odot}$ (from VLTI observations) [31].
Based on these observational data, one can estimate the diameter of the black hole’s shadow per unit mass using the following expression[112, 114, 115]:
@@ III.1 Constraints the black hole parameters with the EHT observational data from $M87^{*}$ and $SgrA^{*}$
The diameter of the shadow can be determined using the formula $d_{\text{sh}}^{\text{theo}}=2R_{\text{sh}}$. Thus, the diameter of the black hole shadow is $d_{\text{sh}}^{\text{Sgr A*}}=(9.5\pm 1.4)M$ for $SgrA^{*}$ and $d_{\text{sh}}^{\text{M87*}}=(11\pm 1.5)M$ for $M87^{*}$
The angular diameters of the black hole shadow in quadratic gravity corresponding to standard photon sphere radius $r_{ph}$ and corresponding to new photon sphere $r^{L}_{ph}$ obtained using the $M87^{*}$ and $SgrA^{*}$ data, are depicted in Fig. 3 and in Fig. 4, respectively.
@@
Table 2: Estimated ranges of the parameters $S_{0}/2M,~{}S_{2}/2M$ and $2Mm_{0},~{}2Mm_{2}$ from the known shadow observables $\Omega(\mu as)$ for $M87^{*}$ and $Sgr~{}A^{*}$ in the context of standard photon sphere radius $r_{ph}$.
@@
Figure 3: The angular diameter $\Omega(\mu as)$ of the black hole shadow correspondence to standard photon sphere radius $r_{ph}/2M$ is presented as a function of the parameters $S_{0},~{}S_{2}$ and $m_{0},~{}m_{2}$ for the $M87^{*}$ (panels (a) and (c)) and $SgrA^{*}$ (panels (b) and (d)) supermassive black holes. The enclosed regions satisfy the 1-$\sigma$ shadow bounds for $M87^{*}$ and $SgrA^{*}$, respectively. In panel (a), the green and blue lines correspond to $\Omega=39.1\,\mu as$ and $\Omega=39.5\,\mu as$, respectively; in panel (b), they correspond to $\Omega=51.4\,\mu as$ and $\Omega=51.8\,\mu as$, respectively; in panel (c), to $\Omega=39.4\,\mu as$ and $\Omega=39.3\,\mu as$, respectively; and in panel (d), to $\Omega=52.2\,\mu as$ and $\Omega=52\,\mu as$, respectively.
Figure 4: The angular diameter $\Omega(\mu as)$ of the black hole shadow correspondence to new photon sphere radius $r^{L}_{ph}/2M$ is presented as a function of the parameters $S_{0},~{}S_{2}$ and $m_{0},~{}m_{2}$ for the $M87^{*}$ (panels (a) and (c)) and $SgrA^{*}$ (panels (b) and (d)) supermassive black holes. The enclosed regions satisfy the 1-$\sigma$ shadow bounds for $M87^{*}$ and $SgrA^{*}$, respectively. In panel (a), the green and blue lines correspond to $\Omega=39.1\,\mu as$ and $\Omega=39.5\,\mu as$, respectively; in panel (b), they correspond to $\Omega=51.4\,\mu as$ and $\Omega=51.8\,\mu as$, respectively; in panel (c), to $\Omega=39.4\,\mu as$ and $\Omega=39.3\,\mu as$, respectively; and in panel (d), to $\Omega=52.2\,\mu as$ and $\Omega=52\,\mu as$, respectively.
Supermassive BH $\Omega(\mu as)$ $S_{0}/2M$ $S_{2}/2M$ $2Mm_{0}$ $2Mm_{2}$
    39.1 $0<S_{0}/2M\leq 0.093$ $0<S_{2}/2M\leq 0.093$     1       1
$M87^{*}$    39.5 $0<S_{0}/2M\leq 0.048$ $0<S_{2}/2M\leq 0.048$     1       1
    39.3      0.01      0.01 $0<2Mm_{0}\leq 0.38$ $0<2Mm_{2}\leq 0.38$
    39.4      0.01      0.01 $0<2Mm_{0}\leq 0.77$ $0<2Mm_{2}\leq 0.77$
    51.4 $0<S_{0}/2M\leq 0.12$ $0<S_{2}/2M\leq 0.12$     1       1
$Sgr~{}A^{*}$    51.8 $0<S_{0}/2M\leq 0.082$ $0<S_{2}/2M\leq 0.082$     1       1
    52.2      0.01      0.01 $0<2Mm_{0}\leq 1.74$ $0<2Mm_{2}\leq 1.74$
    52      0.01      0.01 $0<2Mm_{0}\leq 0.54$ $0<2Mm_{2}\leq 0.54$
Table 3: Estimated ranges of the parameters $S_{0}/2M,~{}S_{2}/2M$ and $2Mm_{0},~{}2Mm_{2}$ from the known shadow observables $\Omega(\mu as)$ for $M87^{*}$ and $Sgr~{}A^{*}$ in the context of new photon sphere radius $r^{L}_{ph}$.
Constraints from $M87^{*}$
@@
The angular diameter of the shadow ($\Omega$) for the $M87^{*}$ black hole, averaged over the range from $32\,\mu as$ to $45\,\mu as$ with a uncertainty region $1\sigma$ from EHT observations, places constraints on the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ of the black hole in quadratic gravity, as illustrated in Figs. 3(a) and 3(c).
@@
Here, the angular diameter of the shadow ($\Omega$) is expressed as a function of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$. In Fig. 3(a), the green and blue lines correspond to $\Omega=39.1\>\mu as$ and $\Omega=39.5\>\mu as$, respectively. Similarly, in Fig. 3(c), the green and blue lines correspond to $\Omega=39.4\>\,\mu as$ and $\Omega=39.3\>\mu as$, respectively. The blue and green curves in Figs. 3(a) and 3(c) align with the shadow size of $M87^{*}$ observed by the EHT. Similar correspondence , the shadow’s angular diameter ($\Omega$) corresponding to new photon sphere radius $r^{L}_{ph}$, is expressed as a function of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see in Fig. 4).
@@
The constraints on the pair of parameters $(S_{0},S_{2})$ corresponding to the radius of the standard photon sphere $r_{ph}$ are given by (see Table 2 ): $0\leq S_{0}/2M\leq 0.09,$ and $0\leq S_{2}/2M\leq 0.09,$ corresponding to $\Omega=39.1\>\mu as$. Similarly, the constraints on the parameter pair $(m_{0},m_{2})$ are: $0\leq 2Mm_{0}\leq 0.78,$ and $0\leq 2Mm_{2}\leq 0.78$, corresponding to $\Omega=39.4\>\mu as$.
The constraints on the parameter pair $(S_{0},S_{2})$ corresponding to new photon sphere radius $r^{L}_{ph}$ are given by (see Table 3 ): $0\leq S_{0}/2M\leq 0.093,$ and $0\leq S_{2}/2M\leq 0.093,$ corresponding to $\Omega=39.1\>\mu as$. Similarly, the constraints on the parameter pair $(m_{0},m_{2})$ are: $0\leq 2Mm_{0}\leq 0.77,$ and $0\leq 2Mm_{2}\leq 0.77$, corresponding to $\Omega=39.4\>\mu as$.
@@
Thus, $M87^{*}$ can be modeled within this constrained parameter space as a black hole in quadratic gravity. The numerous possible parameter points within the parameter spaces $(S_{0},S_{2})$ and $(m_{0},m_{2})$ indicate that the compatibility of the black hole model in quadratic gravity with $M87^{*}$ observations makes it a strong candidate for describing astrophysical black holes.
@@
The angular diameter of the shadow ($\Omega$) for the $SgrA^{*}$ black hole, averaged over the range from $49.5\,\mu as$ to $54.1\,\mu as$ with an uncertainty region $1\sigma$ from EHT observations, places constraints on the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ of the black hole in quadratic gravity, as depicted in Figs. 3(b) and 3(d).
Here, the angular diameter of the shadow ($\Omega$) corresponding to the radius of the standard photon sphere $r_{ph}$, is expressed as a function of the pair of parameters $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see in Fig. 3). In Fig. 3(b), the green and blue lines correspond to $\Omega=51.4\>\mu as$ and $\Omega=51.8\>\mu as$, respectively. Similarly, in Fig. 3(d), the green and blue lines correspond to $\Omega=52.2\>\mu as$ and $\Omega=52\>\mu as$, respectively. The blue and green curves in Figs. 3(b) and 3(d) align with the shadow size of $SgrA^{*}$ observed by the EHT. Similar correspondence , the shadow’s angular diameter ($\Omega$) corresponding to new photon sphere radius $r^{L}_{ph}$, is expressed as a function of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see in Fig. 4).
The constraints on the pair of parameters $(S_{0},S_{2})$ corresponding to the radius of the standard photon sphere $r_{ph}$, are given by (see Table 2): $0\leq S_{0}/2M\leq 0.11,$ and $0\leq S_{2}/2M\leq 0.11,$ corresponding to $\Omega=51.4\>\mu as$. Similarly, the constraints on the parameter pair $(m_{0},m_{2})$ are: $0\leq 2Mm_{0}\leq 1.8$ and $0\leq 2Mm_{2}\leq 1.8$, corresponding to $\Omega=52.2\>\mu as$.
The constraints on the parameter pair $(S_{0},S_{2})$ corresponding to new photon sphere radius $r^{L}_{ph}$, are given by (See Table 3): $0\leq S_{0}/2M\leq 0.12,$ and $0\leq S_{2}/2M\leq 0.12,$ corresponding to $\Omega=51.4\>\mu as$. Similarly, the constraints on the parameter pair $(m_{0},m_{2})$ are: $0\leq 2Mm_{0}\leq 1.74$ and $0\leq 2Mm_{2}\leq 1.74$, corresponding to $\Omega=52.2\>\mu as$.
Thus, $SgrA^{*}$ can be modeled as a black hole in quadratic gravity within this constrained parameter space. The numerous possible parameter points within the parameter spaces $(S_{0},S_{2})$ and $(m_{0},m_{2})$ indicate that the compatibility of the black hole model in quadratic gravity with $SgrA^{*}$ observations makes it a strong candidate for describing astrophysical black holes.
It is found that the parameter ranges obtained for the new photon sphere radius $r^{L}_{ph}$ differ slightly from those obtained for the standard photon sphere radius $r_{ph}$, although finite ranges are obtained in both cases. Thus, the Schwarzschild black hole solution in quadratic gravity emerges as a viable model.